OpenStudy (anonymous):

Solve the equation for solutions in the interval [0, 2pi) sin 2x + sin x = 0

sam (.sam.):

That is minus @FaceYourFears

OpenStudy (anonymous):

^~^ I just gave him a link lol

OpenStudy (anonymous):

I am starting over from @lacypennelll

sam (.sam.):

Right, so @beverly42 are you there?

OpenStudy (linyu):

2sinxcosx+sinx=0 sinx(2cosx+1)=0

OpenStudy (goformit100):

For Solving this Question, I would like to ask you that, do you have the detailed conception of : SetTheory, and Theory of Limiting Numbers

OpenStudy (linyu):

sinx=0, x=k(pi) , where k=0,1,2,........ due to interval [0,2pi) k=0,1 so x=0,pi 2cosx+1=0 2cosx=-1 cosx=-1/2, x=5(pi)/6+2k(pi), x=11(pi)/6+2k(pi) due to interval [0,2pi) k=0 so x=5(pi)/6, 11(pi)/6